A 1-year European call and put options on a non-dividend paying stock has a strike price of 80. You are given: (i) The stock’s price is currently 75.
(ii) The stock’s price will be either 85 or 65 at the end of the year.
(iii) The continuously compounded risk-free rate is 4.5%.

(a) Determine the premium for the call.

(b) Determine the premium for the put.
Answer #1

ANSwey : Strike Price = 80 Current stock price = 75 Stock price can Increase by 85 cor Decrease by 65 Rifk free rate : 4.st.premium of it call option p. Su + C1-p) 0.66875 x 5 + 0.33125 x 0 (uasa 3.34375 2. of & call option 1 year premium after 4- 36) Premium of put option. = pisu) + C1-P) Sd 9-13 with = 0.66875 X0 + 0.33125 x 15 d. 2154 0+ 4.96875 4.96875 A / B gr. premi